OpenStudy (anonymous):

Which should be factored using the Bottoms Up method? (Points : 2) x^2 -7x + 10 x^2 - 121 x^2 + 20x + 100 all should be factored using the bottoms up method. none can be factored using the bottoms up method.

OpenStudy (anonymous):

@iGreen

OpenStudy (welshfella):

lol never heard of the bottoms up method!!

OpenStudy (anonymous):

:(

OpenStudy (xapproachesinfinity):

bottom up method haha what the heck they mean ac method perhaps

OpenStudy (anonymous):

No idea.

OpenStudy (anonymous):

Why does no one know this method?!?!

OpenStudy (xapproachesinfinity):

bottom up method, could be the same as ac just people create nonsense different names to the same thing

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Well could you help me?

OpenStudy (xapproachesinfinity):

they are trying to be artistic ^_-

OpenStudy (anonymous):

oh lol

OpenStudy (xapproachesinfinity):

well clearly the second can be factored using that kind of method rather it is the difference of two squares

OpenStudy (xapproachesinfinity):

so think about the first x^2-7x+10 here a=1 c=10 so we just look for factor of 10 that add up to -7

OpenStudy (xapproachesinfinity):

-2 and -5 are those two numbers so x^2-2x-5x+10=x(x-2)-5(x-2)=(x-2)(x-5)

OpenStudy (xapproachesinfinity):

so we can factor it using ac method

OpenStudy (xapproachesinfinity):

let me see google what it says about this bottom up method of yours haha

OpenStudy (xapproachesinfinity):

well google said it is actually ac method with a>1 so here those given one cannot be applied to this method

OpenStudy (xapproachesinfinity):

eh nonsense about factoring really! answer to your question is no they cannot

OpenStudy (anonymous):

thanks man :) i really appericate it. Could you help with a few more?

OpenStudy (xapproachesinfinity):

clear! don't see why this question is asked there a lack of what for here

OpenStudy (xapproachesinfinity):

go ahead and ask i will leave soon

OpenStudy (anonymous):

Find the 2 solutions for 6x^2 + 11x +3=0 Find the 2 solutions for 10x^2-7x+1=0

OpenStudy (xapproachesinfinity):

use the ac method again take 6x3=18 factors of 18 that add up to -11 so 9 and 2 your equation can be rewrtitten as 6x^2+2x+9x+3=0 2x(3x+1)+3(3x+1)=0 (3x+1)(2x+3)=0 3x+1=0 or 2x+3=0 x=-1/3 or x=-3/2

OpenStudy (xapproachesinfinity):

you will have to do repeat the same steps with the other one

OpenStudy (xapproachesinfinity):

try and let's how you did

OpenStudy (xapproachesinfinity):

so... any attempt?

OpenStudy (anonymous):

sorry. I had to do something

OpenStudy (xapproachesinfinity):

ok once you done doing that then try to do this